OpenStudy (yusufm):

im trying to simplify 2sin2x−2sinx+2∗3√cosx−√3=0 can someone give me a hint as to where to start?

OpenStudy (kky10997):

can you write equation again

OpenStudy (yusufm):

m trying to simplify 2sin2x−2sinx+2∗√3cosx−√2=0 can someone give me a hint as to where to start?

OpenStudy (kky10997):

yeah

OpenStudy (kky10997):

1st apply sin2x=2sinxcosx

OpenStudy (kky10997):

then divide both side by cosx

OpenStudy (yusufm):

ok

OpenStudy (kky10997):

so now you get

OpenStudy (yusufm):

[2sinxcosx-2sinx+2*\sqrt{3}cosx-\sqrt{3}\] then divide by cos [2sinx-2*(sinx/cosx)+2\sqrt{3}-(\sqrt{3}/\cos)\]

OpenStudy (yusufm):

where did you go? I can't figure it out. Please help